# Multi-Virtual Knot Theory

### v1

@@ 5.1 Generalized Biquandles
where is the flat diagram with four virtual crossings shown in the Figure.
Note that and We have already proved that the flat state is non trivial under virtual equivalence.
@@ 7 Parity Bracket Polynomials for Multiple Virtual Knot Theory
a number of non-trivial virtual knots overlying it. The free knots can be studied for their own sake. In order to consider a multi-virtual analog for free knots, one can add chords to chord diagrams that correspond to different virtual crossing types, or one can use diagrams as we do in this paper with classical crossings replaced by flat crossings, and add the -move to the equivalence relation where the -move allows the clasical and any virtual crossing that are adjacent to one another to undergo and interchange. This generalization will be pursued separately from the present paper.
@@ 8 Generalized Arrow Polynomial
In working with multiple virtuals there are many more reduced states as shown at the bottom of the figure. These consist in loops with zig-zags that are bound to one another due to flat linkage of the multiple virtual crossings. Thus, the generalized arrow polynomial is a summation over the equivalence classes of its states, with polynomial coefficients.
See Figure 70 for an example calculation. In this example we show how flat linked virtual states can occur in the detection of a multiple virtual knot.
In Figures 71 and 72 we illustrate two arrow polynomial calculations. In the first figure we have a link with one type of virtual crossing and find that it has
@@ 9 Multiple Welded Knot Theory
Welded knot theory is based on allowing the forbidden move but not allowing the move See Figure 74 and 75.
@@ 10 Rotational Multiple Virtuals
Note that for a finite number of strands the group in involves the classical braid group and as many copies of the symmetric group as there are

### v2

@@ 5.1 Generalized Biquandles
Biquandles are a generalization of quandles where two operations occur at each crossing. We will use exponential notation to describe these operations. We will write
and
It is understood that the “bar” in is strictly notational and does not designate a separate operation on
@@
where is the diagram with four virtual crossings shown in the Figure.
Note that and We have already proved that the state is non trivial under virtual equivalence.
@@ 7 Parity Bracket Polynomials for Multiple Virtual Knot Theory
a number of non-trivial virtual knots overlying it. The free knots can be studied for their own sake. In order to consider a multi-virtual analog for free knots, one can add chords to chord diagrams that correspond to different virtual crossing types, or one can use diagrams as we do in this paper with classical crossings replaced by immersion crossings, and add the -move to the equivalence relation where the -move allows the clasical and any virtual crossing that are adjacent to one another to undergo and interchange. This generalization will be pursued separately from the present paper.
@@ 8 Generalized Arrow Polynomial
In working with multiple virtuals there are many more reduced states as shown at the bottom of the figure. These consist in loops with zig-zags that are bound to one another due to virtual linkage of the multiple virtual crossings. Thus, the generalized arrow polynomial is a summation over the equivalence classes of its states, with polynomial coefficients.
See Figure 74 for an example calculation. In this example we show how linked virtual states can occur in the detection of a multiple virtual knot.
In Figures 75 and 76 we illustrate two arrow polynomial calculations. In the first figure we have a link with one type of virtual crossing and find that it has
@@ 9 Multiple Welded Knot Theory
Welded knot theory is based on allowing the forbidden move but not allowing the move See Figure 78 and 79.
@@ 10 Rotational Multiple Virtuals
Note that for a finite number of strands the group involves the classical braid group and as many copies of the symmetric group as there are